OpenStudy (anonymous):

y – 0.6 = –1.8 A. y = –2.4 B. y = –1.2 C. y = 1.2 D. y = 2.4

rishavraj (rishavraj):

add 0.6 on both sides......

OpenStudy (anonymous):

153.0

OpenStudy (lolacole12):

Let's solve your equation step-by-step. y−0.6=−1.8 Step 1: Add 0.6 to both sides. y−0.6+0.6=−1.8+0.6 y=−1.2 Answer: y=−1.2

OpenStudy (lolacole12):

Understand

OpenStudy (anonymous):

b

OpenStudy (anonymous):

shure??

OpenStudy (anonymous):

yeah

OpenStudy (lolacole12):

Huzaayfa i already explained it seems like you just copied

OpenStudy (anonymous):

i didnt

OpenStudy (mathmale):

ConfusedBoy: If you need to ask "shure," (actually it's "sure?"), that indicates you need to get the practice and build confidence in your ability to solve these problems yourself.

OpenStudy (lolacole12):

me but good job

OpenStudy (anonymous):

who me

OpenStudy (mathmale):

y – 0.6 = –1.8 confusedboy: How would you move that -1.6 to the right side, to isolate y on the left side?

OpenStudy (lolacole12):

@Huzayfa its ok i belive you good job you get medal

OpenStudy (anonymous):

oh i am from poland so i have an accent

OpenStudy (anonymous):

but ok

OpenStudy (mathmale):

@huzayfa and @Iolacole12: enough, please! Your job is to help Confusedboy learn how to do these problems with your guidance and then by himself. Kindly do not work out the entire problem and present the answer without his involvement. y – 0.6 = –1.8 Confusedboy: How would we isolate y (by moving the -0.6 term to the right side of the equation)?

OpenStudy (anonymous):

that is how we say shure but ok sure

OpenStudy (anonymous):

thanks lolacole12

OpenStudy (anonymous):

its ok but real people that dont make fun of my name helped me